OpenStudy (anonymous):

Need help writing an equation: A TANK IS FILLING WITH WATER FROM A NATURAL SPRING. TWO DAYS AGO THE WATER WAS 10 FEET DEPTH, AND YESTERDAY THE WATER WAS 12 FEET. ASSUME THAT THE WATER DEPTH CONTINUES TO RISE AT THIS RATE AFTER TODAY. I simply need to know what the expression would be for this equation since anything Im putting is wrong. really doesnt make sense to me

OpenStudy (anonymous):

What is the difference in depth each day?

OpenStudy (anonymous):

What is the initial depth in the tank? i.e. Day 1

OpenStudy (anonymous):

the difference each day is 2 feet

OpenStudy (anonymous):

someone already asked this question but it was unclear what the final answer was :/

OpenStudy (anonymous):

I'd say it would be Depth = 2ft times (#days) + 10ft f(x) = 2x + f(0) where f(x) is total depth, x is days, and f(0) is initial depth

OpenStudy (anonymous):

sorry............

OpenStudy (anonymous):

its cool

OpenStudy (anonymous):

@ChmE explained the best way it is

OpenStudy (compassionate):

HI, SO I SEE YOU'RE LOOKING TO SOLVE AN EQUATION. WELL, BASICALLY, YOU SEE IT HAS AN INCREASE OF 2. SO YOU CAN PUT F(X) = 2X + 10, YOU SEE? IT RISES BY TWO FEET, SO WE MULTIPLY THE DAYS BY TWO, WE ADD ANY INITIAL DEPTH GIVEN, AND ADD TEN BECAUSE THAT IS WHERE WE STARTED. NOW FOR ANY OTHER QUESTION YOU CAN SAY F(X) = 2X + F(0), AND PLUG IN THE INITAL DEPTH TO F(0) I HOPE THIS HELPED, IN CAPS LOCK AND COMPASSION, COMPASSIOANTE

OpenStudy (anonymous):

guys im going to post a snip image of the problem, so you can have a better reference, because anything I put in here does not work, I literally give up its been 2 hours ive been trying to solve this stupid problem

OpenStudy (phi):

let days be "x", and depth be "y" "2 days ago the water was 10 feet deep" can be translated to the point (-2,10) (we will measure days from NOW, so 2 days ago is -2) "1 day ago the wage was 12 feet deep" becomes (-1,12) now we have two points. Find the equation of a line using 2 points. First, the slope is change in y divided by change in x: (12-10)/ (-1 - (-2))= 2/1=2 now use the point-slope formula y - y0= m(x-x0) y-12 = 2(x - (-1)) which simplifies y-12 = 2(x +1) --> y -12 = 2x + 2 --> y=2x+14 now that you have your formula you can answer the questions.

OpenStudy (phi):

however, exactly what your computer wants to see for the equation is unclear. they may want you to use "d" instead of x? they may want to just see 2d+14 (with no y?)

OpenStudy (anonymous):

omg thank you phi , it was 2d+14 lol, u deserve a 1000 medals